# Computer-Assisted Proofs of Gap Solitons in Bose-Einstein Condensates

### v2

@@ Abstract
We provide a framework for turning a numerical simulation of a gap soliton in the one-dimensional Gross-Pitaevskii equation into a formal mathematical proof of its existence. These nonlinear localized solutions play a central role in understanding Bose-Einstein condensates (BECs). We reformulate the problem of proving their existence as the search for homoclinic orbits in a dynamical system. We then apply computer-assisted proof techniques to obtain verifiable conditions under which a numerically approximated trajectory corresponds to a true homoclinic orbit. This work also presents the first examples of computer-assisted proofs of gap solitons in the Gross-Pitaevskii equation on non-perturbative parameter regimes.
@@ 1 Introduction
A Bose-Einstein condensate (BEC) is a state of matter that forms when a collection of particles cools down to temperatures near absolute zero, causing them to lose their individual identities and merge into a single wave. The first experimental realization of BECs at ultra-cold temperatures earned the 2001 Nobel Prize in Physics, and since then, Bose-Einstein condensates have provided a platform for exploring quantum mechanics on large scales, with applications in precision measurements, quantum computing, and the modeling of complex systems such as superfluidity and optical lattices. For a comprehensive review of both experimental and theoretical developments, see [1].
Beyond their physical significance, the study of Bose-Einstein condensates offers fertile ground for advancing theoretical methods in nonlinear dynamics and partial differential equations. In this paper, we study the dynamics of a BEC using the time-dependent Gross-Pitaevskii equation, which models the BEC’s evolution in one spatial dimension:
@@
Here, $\psi(t,x)\in\mathbb{C}$ denotes the dimensionless wave function, $|\psi|^{2}$ represents the BEC density, and $V(x)$ is the external potential created by the optical lattice along the spatial domain $x\in\mathbb{R}$. The book [2] offers a comprehensive treatment of the Gross-Pitaevskii equation and nonlinear wave dynamics, combining experimental perspectives with numerical studies of Bose-Einstein condensates. It also includes an extensive bibliography covering many areas of the field.
A fundamental aspect of Bose-Einstein condensate analysis is the study of standing wave solutions to the Gross-Pitaevskii equation. These solutions take the form $\psi(t,x)=e^{-iat}u(x)$, where $u(x)$ is a real-valued function that satisfies the time-independent (GP) equation:
@@
Chapter 19 of [2] surveys the existence of solitons under various potential types, with particular emphasis on periodic potentials. In the present work, we study the specific case $V(x)=b\cos(2x)$, which models the dynamics of a Bose-Einstein condensate in an optical lattice. To understand the setting in which gap solitons arise, we examine the linearization of the GP equation around the trivial solution, called Mathieu’s equation:
[…]
@@
Researchers have frequently used perturbative asymptotics and numerical simulations to investigate soliton solutions. For example, [3] applies asymptotic analytical methods to compute gap solitons in all spectral gaps of a periodic potential, showing that these solitons bifurcate from distinct band edges depending on the sign of $c$. Moreover, [4] introduces a numerical approach for calculating gap solitons in the repulsive case ($c=1$). This study shows that most solutions blow up at certain points on the real axis. Notably, the subset of non-blow-up solutions—which includes all gap solitons—forms a fractal set within the space of initial conditions. This fractal structure makes it possible to identify gap solitons over large regions in the parameter space $(a,b)\in\mathbb{R}^{2}$. While simulations and asymptotic techniques provide valuable insight, they do not yield a rigorous proof of existence. Abstract results, on the other hand, often lack any explicit description of the solution profiles.
@@
The Gross-Pitaevskii equation (2) with parameters $a=1.1025$, $b=0.55125$, and $c=-0.826875$ has a soliton solution $u:\mathbb{R}\to\mathbb{R}$, satisfying
@@
Figure 1: The figure shows a validated soliton solution of the Gross-Pitaevskii (GP) equation with parameters $a=1.1025$, $b=0.55125$, and $c=-0.826875$. It also depicts the main elements of our approach: the solution to the boundary-value problem (blue), the stable manifold (orange), and the even extension of the soliton (black).
[…]
with initial conditions $u_{3}(0)=1$ and $u_{4}(0)=0$. Then, if $u_{1},u_{2}\neq 0$ and the conditions in (3) are satisfied, the first component $u_{1}$ of $U$ is a gap soliton of the GP equation (2). We emphasize that the system (5) is not conservative, though it possesses the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$. This comes from the fact that $(u_{3}(x),u_{4}(x))=(V(x),V^{\prime}(x))$ solves the Hamiltonian system $u_{3}^{\prime}=\partial_{u_{4}}H(u_{3},u_{4})$, $u_{4}^{\prime}=-\partial_{u_{3}}H(u_{3},u_{4})$. Now, define the set
[…]
@@
The study of connecting orbits in dynamical systems is a vast and active area of research, intersecting with diverse mathematical disciplines such as algebraic topology, Morse homology, celestial mechanics, chaos theory, and the calculus of variations. Over the years, a variety of mathematical techniques have been developed to address the theoretical challenges associated with these orbits. For instance, perturbative methods within variational frameworks [6, 7, 8] and non-perturbative techniques [9, 10, 11] have been successfully employed to establish the existence of homoclinic orbits in conservative and Hamiltonian systems. Additionally, the advent of computer-assisted proofs in nonlinear analysis has greatly enriched the study of connecting orbits, employing techniques that leverage the strengths of topology, functional analysis, and scientific computing. Prominent methodologies include topological covering relations [12, 13], the Parameterization Method combined with a functional analytic setup [14, 15, 16, 17], homotopy methods [18], and interpolation-based techniques [19, 20]. These methods have been instrumental in establishing the existence of connecting orbits in the context of ordinary differential equations (ODEs) and continue to play a pivotal role in advancing the field.
@@
Let $X$ and $Y$ be Banach spaces, and let $F:X\to Y$ be a $C^{1}$ map, let $\bar{x}$ be an element of $X$, $A$ a linear injective map from $Y$ to $X$. Let $r^{*}$ be a positive real number and denote by $B(\bar{x},r^{*})$ the closed ball centered at $\bar{x}$ of radius $r^{*}$. Assume there exist nonnegative constant Y, $Z_{1}$ and $Z_{2}$ such that
@@
If $Z_{1}<1$ and $Z_{2}<\frac{(1-Z_{1})^{2}}{2Y}$, then for any $r$ satisfying $\frac{1-Z_{1}-\sqrt{(1-Z_{2})^{2}-2YZ_{2}}}{Z_{2}}\leq r<min\left(\frac{1-Z_{1}}{Z_{2}},r^{*}\right)$, then there exists a unique zero $x^{*}$ of $F$ in the ball $B(\bar{x},r)$.
In our setting, $\bar{x}$ denotes a numerical approximate solution. The quantities $Y$, $Z_{1}$, and $Z_{2}$ correspond to bounds on the norms of elements in certain infinite-dimensional Banach spaces. For any given set of parameters, we derive these bounds in a form that can be rigorously evaluated by a computer. This derivation constitutes the main technical part in our paper.
@@ 2 Computation of a Local Stable Manifold of the Periodic Orbit
In Section 1, we reformulated the problem of proving the existence of soliton solutions as finding a solution to equation (5) that intersects the local stable manifold $W_{loc}^{s}(\gamma)$ associated with the periodic orbit $\gamma$. To explicitly characterize points on this stable manifold, we will employ the Parameterization Method [21, 22, 23], following the framework developed in [30, 24]. The periodic orbit $\gamma$ possesses two trivial Floquet exponents: one arising from the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$, discussed in Section 1, and another due to the shift invariance of the periodic orbit. Consequently, $\gamma$ admits at most two nontrivial Floquet exponents. For the remainder of this work, we assume that $\dim W^{u}(\gamma)=\dim W^{s}(\gamma)=1$ and denote the stable Floquet exponent of $\gamma$ by $\lambda<0$. Denote by $v:S^{1}\to\mathbb{R}^{4}$ the associated stable bundle, that is a solution of
$\dot{v}+\lambda v=Dg(\gamma(\theta))v.$ (11)
We refer to the image of $v$ as the stable tangent bundle attached to the periodic orbit $\gamma$.
The Parameterization Method allows us to compute a parameterization $W:S^{1}\times[-1,1]\to\mathbb{R}^{4}$ of $W_{loc}^{s}(\gamma)$ by solving the following partial differential equation
@@
for a given exponential weight $\nu\geq 1$. Given two sequences of complex numbers $u_{1}=\{(u_{1})_{m}\}_{m\in\mathbb{Z}},u_{2}=\{(u_{2})_{m}\}_{m\in\mathbb{Z}}\in S_{\scriptscriptstyle{\mathsf{F}}}$, denote their discrete convolution given component-wise by
@@
This gives rise to a product $*_{\scriptscriptstyle{\mathsf{F}}}:S_{\scriptscriptstyle{\mathsf{F}}}\times S_{\scriptscriptstyle{\mathsf{F}}}\to S_{\scriptscriptstyle{\mathsf{F}}}$. To represent the Taylor-Fourier coefficients of the parameterization $W$, we consider the sequence space $S_{\scriptscriptstyle{\mathsf{TF}}}$ defined by
@@
Given a Taylor-Fourier sequence $p\in S_{\scriptscriptstyle{\mathsf{TF}}}$, and given $n\geq 0$, we denote by $p_{n}\in S_{\scriptscriptstyle{\mathsf{F}}}$ the Fourier sequence $p_{n}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(p_{n,m})_{m\in\mathbb{Z}}$. Now, given $p,q\in S_{\scriptscriptstyle{\mathsf{TF}}}$ we define their Taylor-Fourier Cauchy product $*_{\scriptscriptstyle{\mathsf{TF}}}:S_{\scriptscriptstyle{\mathsf{TF}}}\times S_{\scriptscriptstyle{\mathsf{TF}}}\to S_{\scriptscriptstyle{\mathsf{TF}}}$ as follows
@@
Since $\lambda<0$, the linear system admits a unique solution, which is the zero vector for all $m\in\mathbb{Z}$. Consequently, $v_{m}^{(3)}=v_{m}^{(4)}=0$ for all $m\in\mathbb{Z}$. Thus, it remains to rigorously compute the coefficients of $v^{(1)}$ and $v^{(2)}$, which we carry out in Section 2.1. Assuming that this is done, from the Fourier-Taylor expansion (19), we then get that $w_{1,m}^{(j)}=v_{m}^{(j)}$ for all $m\in\mathbb{Z}$ and $j=1,2,3,4$.
@@
where the Taylor-Fourier Cauchy product $*_{\scriptscriptstyle{\mathsf{TF}}}$ is given in (18). By applying a similar argument to the one used to establish that $v^{(3)}=v^{(4)}=0$, we conclude that $w_{n,m}^{(3)}=w_{n,m}^{(4)}=0$ for all $n\geq 2$ and $m\in\mathbb{Z}$. Thus, the remaining task in parameterizing $W_{loc}^{s}(\gamma)$ reduces to rigorously enclosing the coefficients $w_{n,m}^{(1)}$ and $w_{n,m}^{(2)}$ for all $n\geq 2$ and $m\in\mathbb{Z}$. In the remainder of this section, we develop a general computer-assisted framework to prove the existence of solutions to the first two equations of the sequence equations (23) and (25) close to approximate solutions. For each problem, we construct a validation map $F$, whose zeros correspond to the desired solutions. The existence of these solutions is then established using the (Newton-Kantorovich) Theorem 2.

### v3

@@ Abstract
We provide a framework for turning a numerical simulation of a gap soliton in the one-dimensional Gross-Pitaevskii equation into a rigorous mathematical proof of its existence. These nonlinear localized solutions play a central role in the study of Bose-Einstein condensates (BECs). We reformulate the problem of proving their existence as the search for homoclinic orbits in a dynamical system. We then apply computer-assisted proof techniques to obtain verifiable conditions under which a numerically approximated trajectory corresponds to a true homoclinic orbit. This work also presents the first examples of computer-assisted proofs of gap solitons in the Gross-Pitaevskii equation on non-perturbative parameter regimes.
@@ 1 Introduction
A Bose–Einstein condensate (BEC) is a state of matter that occurs when a collection of particles cools down to temperatures near absolute zero, causing them to lose their individual identities and behave as a single wave. The first experimental realization of BECs at ultra-cold temperatures earned the 2001 Nobel Prize in Physics, and since then, Bose-Einstein condensates have provided a platform for exploring quantum mechanics on large scales, with applications in precision measurements, quantum computing, and the modeling of complex systems such as superfluidity and optical lattices. For a comprehensive review of both experimental and theoretical developments on Bose–Einstein condensates, see [22].
Beyond their physical significance, the study of Bose–Einstein condensates offers fertile ground for advancing theoretical methods in nonlinear dynamics and partial differential equations. In this paper, we study the dynamics of a BEC using the time-dependent Gross–Pitaevskii equation, which models the BEC’s evolution in one spatial dimension:
@@
Here, $\psi(t,x)\in\mathbb{C}$ denotes the dimensionless wave function, $|\psi|^{2}$ represents the BEC density, and $V(x)$ is the external potential created by the optical lattice along the spatial domain $x\in\mathbb{R}$. The book [19] provides a detailed treatment of the Gross–Pitaevskii equation and nonlinear wave dynamics, combining experimental perspectives with numerical studies of Bose–Einstein condensates. It also includes an extensive bibliography covering many areas of the field.
A fundamental aspect of Bose-Einstein condensate analysis is the study of standing wave solutions to the Gross–Pitaevskii equation. These solutions take the form $\psi(t,x)=e^{-iat}u(x)$, where $u(x)$ is a real-valued function that satisfies the time-independent (GP) equation:
@@
Chapter 19 of [19] surveys the existence of solitons under various potential types, with particular emphasis on periodic potentials. In the present work, we study the specific case $V(x)=b\cos(2x)$, which models the dynamics of a Bose-Einstein condensate in an optical lattice. To understand the setting in which gap solitons arise, we consider the linearization of the Gross–Pitaevskii equation around the trivial solution, known as Mathieu’s equation:
$Lu\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,\left(\partial_{x}^{2}+a-V(x)\right)u=\left(\partial_{x}^{2}+a-b\cos(2x)\right)u=0.$ (4)
@@
Researchers have frequently used perturbative asymptotics and numerical simulations to investigate soliton solutions. For example, [27] applies asymptotic analytical methods to compute gap solitons in all spectral gaps of a periodic potential, showing that these solitons bifurcate from distinct band edges depending on the sign of $c$. Moreover, [2] introduces a numerical approach for calculating gap solitons in the repulsive case ($c=1$). The study shows that the subset of non-blow-up solutions, which includes all gap solitons, has a fractal structure in the space of initial conditions. This fractal structure makes it possible to identify gap solitons over large regions in the parameter space $(a,b)\in\mathbb{R}^{2}$. However, while simulations and asymptotic techniques provide valuable insight, they do not provide a rigorous proof of existence. Abstract results, on the other hand, often lack any explicit description of the solution profiles.
@@
The Gross-Pitaevskii equation (2) with parameters $a=1.1025$, $b=0.55125$, and c=0.826875c=-0.826875 has a soliton solution $u:\mathbb{R}\to\mathbb{R}$, satisfying
@@
Figure 1: The figure shows a validated soliton solution of the Gross-Pitaevskii (GP) equation with parameters $a=1.1025$, $b=0.55125$, and c=0.826875c=-0.826875. It also depicts the main elements of our approach: the solution to the boundary-value problem (blue), the stable manifold (orange), and the even extension of the soliton (black).
To describe our approach, we begin by reformulating the problem using a standard dynamical systems framework. In this setting, finding a gap soliton becomes the search for a connecting orbit between invariant sets. More precisely, denote $u_{1}\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,u$, $u_{2}\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,u^{\prime}$, $u_{3}\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,V(x)=\cos(2x)$ (the periodic potential) and $u_{4}\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,u_{3}^{\prime}=-2\sin(2x)$. Note that $u_{3}$ solves $u_{3}^{\prime\prime}=-4u_{3}$ with initial conditions $(u_{3}(0),u_{3}^{\prime}(0))=(1,0)$. This formulation allows us to transform equation (2) into a autonomous polynomial vector field. Indeed, assume that $U\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,(u_{1},u_{2},u_{3},u_{4})\in\mathbb{R}^{4}$ is a solution to the four-dimensional autonomous system
$\frac{dU}{dx}=g(U)\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,\begin{pmatrix}u_{2}\\ -au_{1}+bu_{3}u_{1}+cu_{1}^{3}\\ u_{4}\\ -4u_{3}\end{pmatrix}$ (5)
with initial conditions $u_{3}(0)=1$ and $u_{4}(0)=0$. Then, if $u_{1},u_{2}\neq 0$ and the conditions in (3) are satisfied, the first component $u_{1}$ of $U$ is a gap soliton of the GP equation (2).
We emphasize that the system (5) is not conservative, though it possesses the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$. This comes from the fact that $(u_{3}(x),u_{4}(x))=(V(x),V^{\prime}(x))$ solves the Hamiltonian system $u_{3}^{\prime}=\partial_{u_{4}}H(u_{3},u_{4})$, $u_{4}^{\prime}=-\partial_{u_{3}}H(u_{3},u_{4})$. Our method can be extended to periodic solutions $V(x)$, where $(u_{3},u_{4})=(V,V^{\prime})$ satisfies a general polynomial Hamiltonian system with Hamiltonian $H(u_{3},u_{4})$. For analytic Hamiltonians that are not polynomial, recently developed techniques based on the discrete Fourier transform could be adapted to our setting to achieve similar results (see, e.g., [14, 35]). Thus, although we have implemented our procedure specifically for the prototype case of the cosine potential, it is applicable to a broad class of other periodic potentials.
Now, define the set
$\gamma\,\overset{\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}{=}\,\{(0,0,\cos(2x),-2\sin(2x)):x\in[0,\pi)\}\subset\mathbb{R}^{4},$ (6)
@@
The study of connecting orbits in dynamical systems is a vast and active area of research that intersects with diverse mathematical disciplines such as algebraic topology, Morse homology, celestial mechanics, chaos theory, and the calculus of variations. Over the years, a variety of mathematical techniques have been developed to address the theoretical challenges associated with these orbits. For instance, perturbative methods within variational frameworks [5, 7, 3] and non-perturbative techniques [28, 18, 13] have been employed to prove the existence of homoclinic orbits in conservative and Hamiltonian systems. More recently, the advent of computer-assisted proofs in nonlinear analysis has greatly enriched the study of connecting orbits, employing techniques that combine the strengths of topology, functional analysis, and scientific computing. Prominent methodologies include topological covering relations [37, 36], the Parameterization Method in a functional-analytic setting [34, 32, 30, 24], homotopy methods [6], and interpolation-based techniques [17, 26]. These methods have been instrumental in proving the existence of connecting orbits in the context of ordinary differential equations (ODEs) and continue to play an important role in advancing the field. In this paper, we contribute to this line of work.
@@
Let $X$ and $Z$ be Banach spaces, and let $F:X\to Z$ be a $C^{1}$ map, let $\bar{x}$ be an element of $X$, AA a linear injective map from $Z$ to $X$. Let $r^{*}$ be a positive real number and denote by $B(\bar{x},r^{*})$ the closed ball centered at $\bar{x}$ of radius $r^{*}$. Assume there exist nonnegative constant Y, $Z_{1}$ and $Z_{2}$ such that
@@
If $Z_{1}<1$ and $Z_{2}<\frac{(1-Z_{1})^{2}}{2Y}$, then for any $r$ satisfying $\frac{1-Z_{1}-\sqrt{(1-Z_{1})^{2}-2YZ_{2}}}{Z_{2}}\leq r<min\left(\frac{1-Z_{1}}{Z_{2}},r^{*}\right)$, then there exists a unique zero $x^{*}$ of $F$ in the ball $B(\bar{x},r)$.
A proof of this theorem can be found in [31]. In our setting, $\bar{x}$ denotes a numerical approximate solution. The quantities $Y$, $Z_{1}$, and $Z_{2}$ correspond to bounds on the norms of elements in certain infinite-dimensional Banach spaces. For any given set of parameters, we derive these bounds in a form that can be rigorously evaluated by a computer. This derivation constitutes the main technical part in our paper.
@@ 2 Computation of a Local Stable Manifold of the Periodic Orbit
In Section 1, we reformulated the problem of proving the existence of soliton solutions as finding a solution to equation (5) that intersects the local stable manifold $W_{loc}^{s}(\gamma)$ associated with the periodic orbit $\gamma$. To explicitly characterize points on this stable manifold, we will employ the Parameterization Method [8, 9, 10], following the framework developed in [11, 12]. The periodic orbit $\gamma$ possesses two trivial Floquet exponents: one arising from the conserved quantity $H=\frac{1}{2}(u_{4}^{2}+4u_{3}^{2})$, discussed in Section 1, and another due to the shift invariance of the periodic orbit. Consequently, $\gamma$ admits at most two nontrivial Floquet exponents. For the remainder of this work, we assume that $\dim W^{u}(\gamma)=\dim W^{s}(\gamma)=1$ and denote the stable Floquet exponent of $\gamma$ by $\lambda<0$. To compute the stable bundle $v$ associated with the periodic solution $\gamma$, we linearize the vector field $g$ in (5) around $\gamma$, yielding the linearized equation
$\dot{u}(\theta)=Dg(\gamma(\theta))\,u(\theta),$ (11)
where the matrix $Dg(\gamma(\theta))$ is $\pi$-periodic. By Floquet theory, solutions of this linear system can be expressed as
$u(\theta)=e^{\lambda\theta}v(\theta),$ (12)
with $\lambda\in\mathbb{R}$ the Floquet exponent and $v(\theta)$ a $2\pi$-periodic function. The period of $v(\theta)$ is doubled relative to that of the original solution to account for the possible non-orientability of the stable bundle. Plugging the expression (12) in (11) yields the functional equation
$\dot{v}+\lambda v=Dg(\gamma(\theta))v,$ (13)
where the solution $v:S^{1}\to\mathbb{R}^{4}$ denotes the associated stable bundle. We refer to the image of $v$ as the stable tangent bundle attached to the periodic orbit $\gamma$.
Once the solution $v$ of (13) is obtained, the Parameterization Method allows us to compute a parameterization $W:S^{1}\times[-1,1]\to\mathbb{R}^{4}$ of $W_{loc}^{s}(\gamma)$ by solving the following partial differential equation
@@
for a given exponential weight $\nu\geq 1$.
Given two sequences of complex numbers $u_{1}=\{(u_{1})_{m}\}_{m\in\mathbb{Z}},u_{2}=\{(u_{2})_{m}\}_{m\in\mathbb{Z}}\in S_{\scriptscriptstyle{\mathsf{F}}}$, denote their discrete convolution given component-wise by
@@
This gives rise to a product $_{\scriptscriptstyle{\mathsf{F}}}:S_{\scriptscriptstyle{\mathsf{F}}}\times S_{\scriptscriptstyle{\mathsf{F}}}\to S_{\scriptscriptstyle{\mathsf{F}}}$. To represent the Taylor-Fourier coefficients of the parameterization $W$, we consider the sequence space $S_{\scriptscriptstyle{\mathsf{TF}}}$ defined by
@@
Given a Taylor-Fourier sequence $p\in S_{\scriptscriptstyle{\mathsf{TF}}}$, and given $n\geq 0$, we denote by $p_{n}\in S_{\scriptscriptstyle{\mathsf{F}}}$ the Fourier sequence $p_{n}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(p_{n,m})_{m\in\mathbb{Z}}$. Now, given $p,q\in S_{\scriptscriptstyle{\mathsf{TF}}}$ we define their Taylor-Fourier Cauchy product $_{\scriptscriptstyle{\mathsf{TF}}}:S_{\scriptscriptstyle{\mathsf{TF}}}\times S_{\scriptscriptstyle{\mathsf{TF}}}\to S_{\scriptscriptstyle{\mathsf{TF}}}$ as follows
@@
Since $\lambda<0$, the linear system admits a unique solution, which is the zero vector for all $m\in\mathbb{Z}$. Consequently, $v_{m}^{(3)}=v_{m}^{(4)}=0$ for all $m\in\mathbb{Z}$. Thus, it remains to rigorously compute the coefficients of $v^{(1)}$ and $v^{(2)}$, which we carry out in Section 2.1. After obtaining a proof in the sequence space $\scriptscriptstyle{\mathsf{F}}$, we show that the Fourier coefficients $v_{m}$ of the solution have the property that the negative modes are the complex conjugates of the corresponding positive modes, so the associated Fourier series represents a real-valued periodic function. Assuming that this is done, from the Fourier-Taylor expansion (21), we then get that $w_{1,m}^{(j)}=v_{m}^{(j)}$ for all $m\in\mathbb{Z}$ and $j=1,2,3,4$.
@@
where the Taylor-Fourier Cauchy product $_{\scriptscriptstyle{\mathsf{TF}}}$ is given in (20). By applying a similar argument to the one used to establish that $v^{(3)}=v^{(4)}=0$, we conclude that $w_{n,m}^{(3)}=w_{n,m}^{(4)}=0$ for all $n\geq 2$ and $m\in\mathbb{Z}$. Thus, the remaining task in parameterizing $W_{loc}^{s}(\gamma)$ reduces to rigorously enclosing the coefficients $w_{n,m}^{(1)}$ and $w_{n,m}^{(2)}$ for all $n\geq 2$ and $m\in\mathbb{Z}$. In the remainder of this section, we develop a general computer-assisted framework to prove the existence of solutions to the first two equations of the sequence equations (25) and (27) close to approximate solutions. For each problem, we construct a validation map $F$, whose zeros correspond to the desired solutions. We then apply the Newton–Kantorovich Theorem 2 to prove the existence of these solutions.